# Hybrid Adaptive Modeling in Process Monitoring: Leveraging Sequence Encoders and Physics-Informed Neural Networks

### v1

@@ 1 Introduction
The rapid growth of research in physics-informed machine learning [3] (PIML) has led to innovative methods that integrate physical knowledge into data-driven models. These approaches benefit from advancements in both machine learning algorithms and hardware acceleration, making them increasingly accessible. Unlike conventional neural networks, physics-informed models incorporate governing physical laws into the learning process, improving their ability to generalize beyond available supervised data. A notable example is Physics-Informed Neural Networks (PINNs), which leverage both physics-based constraints and observed data to solve Partial and Ordinary Differential Equations (PDEs/ODEs). PINNs have gained considerable interest due to their successful applications in fields such as engineering, or fluid dynamics, as well as their ability to handle multi-fidelity data sources [4]. They have also been benchmarked against Finite Element Methods (FEM) [5], demonstrating very good performance. Furthermore, PINNs have shown robustness in solving inverse problems, effectively handling noisy data to infer accurate underlying dynamics. This ability to explore physics through data-driven methods has become a highly active research area.
@@ 2 Identification methods and PINN
[…]
@@ 4 Architecture of the combination of Sequence Encoder and Physics Informer Neural Networks
Figure 6: Overview of the PINN-SE model that combines two Sequences Encoders (Deep Sets or Set Transformer) with PINN and identification blocks, along with the training logic of the model. In this example architecture, we use two Sequence Encoders: the first to encode sensor data $s_{i}$, and the second to encode boundary conditions ($BCs$). Each SE generates a feature vector, which are then concatenated into a global feature vector $\mathbf{F}_{\text{vect}}$ and used by both the PINN and the identification blocks.
@@ 5.2.1 Data generation
$\sigma$ $=\alpha\cdot\frac{\sqrt{\sum_{i=1}^{N}I_{\text{exact}}^{j}(t_{i})^{2}}}{N},\quad\alpha\equiv\text{ratio}\in[0\%,~{}10\%],\quad j\in\{x(t),~{}y(t),~{}z(t)\}$
@@ 5.2.2 Model architecture and training
The model is trained using the expanding window method (Example in Fig. 10). This approach leverages the advantages of Deep sets, that can encode variable sequence lengths and map them to dynamic behaviors. This means the model is trained to integrate new time snapshot data to update its predictions. In this application, the model is trained using 362 datasets selected randomly (using Sobol sampling), the rest is used for validation and testing. The model architecture is given in Fig. 9. The model input, $X_{DS}$, represents time and corresponding state variable snapshots. We use a time step of $\Delta T_{DS}=0.5~{}s$ for the Deep sets input. The input tensor is $X_{DS}\in\mathbb{R}^{b\times N\times d}$, where $b$ is the batch size, $N\in[10,40]$ is the number of snapshot, and $d=4$ is the input dimension (time, $x$, $y$, and $z$). This input is projected to a higher-dimensional representation $\Phi\in\mathbb{R}^{b\times N\times d_{f.v}}$. We then apply an aggregation function to compute the feature vector $\mathbf{F}_{\text{vect}}\in\mathbb{R}^{b\times 1\times d_{f.v}}$. The choice of aggregation function is crucial and must be made carefully to ensure generalization. In this application, we used the following aggregation Eq. 13:
.$\left\|.\right\|_{2}^{2}\equiv\sum_{i=1}^{N_{time}}\frac{\Phi_{i}^{2}}{N}.$ (13)
@@
Figure 9: The PINN-SE model used for the first application (Rössler dynamics) operates as follows: the Deep set takes as input the $time$ and state variables $x$, $y$, and $z$, given as $X_{DS}\in\mathbb{R}^{\text{b}\times N\times 4}$, and predicts feature vector $\mathbf{F}_{\text{vect}}\in\mathbb{R}^{b\times 1\times d_{f.v}}$. This feature vector is then used directly by the identification blocks (MLPs) to predict the parameter $c$. The feature vector is also concatenated with the time snapshots for which we want to predict the state variables $x$, $y$, and $z$ using the PINN blocks. The dynamics loss (Eq. 15) is calculated using automatic differentiation.
@@
Where $\mathcal{L}_{\text{sup}}$ is the supervised loss, and $I(t)=[x(t),~{}y(t),~{}z(t)]$ represents the labeled noisy points used for supervision loss. We employed sparse input sampling with $\Delta T_{\text{sup}}=0.5$ for the supervised loss. The model prediction, $\widetilde{I(t)}$, is computed as shown in Eq. 16:
@@
Deep set 80 4 $\text{gelu}(X)$
Pooling - .$\left\|.\right\|_{2}^{2}$
Post-Pooling 80 3 $\text{gelu}(X)$
@@
Identification 80 3 $\text{gelu}(X)$
@@
12:     if Validation error of parameters $<$ threshold then
@@ 5.2.3 Results and analysis
We also tested the ability to use cross validation between the identification block and the PINN to provide a confidence level for the model’s prediction. To do so, we calculated the residual of the third equation of the Rössler dynamics (eq. 9, $\mathcal{L}_{\text{ODE}}$), by computing $\dot{z}$ using automatic differentiation and the Rössler dynamics corresponding to $b+\widetilde{z}(\widetilde{x}-\widetilde{c})$, where $\widetilde{c}$ is the prediction of the identification model. This was used to calculate the ODE loss $\mathcal{L}_{\text{ODE}_{z}}$ and $\mathbf{MAE_{z}}$. In Fig. 18, we present the correlation between $\mathcal{L}_{\text{ODE}_{z}}=\left\|\dot{z}-b+\widetilde{z}(\widetilde{x}-\widetilde{c})\right\|_{1}$ and the $\mathbf{MAE_{z}}$ loss for the state variable $z$ across all the test datasets.
@@ 5.3.1 Data generation and preparation
[…]
@@ 5.3.2 Model architecture and training process
$\textbf{X}_{DS}=\{t,~{}p_{1},~{}\cdots,~{}p_{i},~{}\cdots,~{}p_{12}\}\in\mathbb{R}^{b\times N_{DS}\times 13}\quad i\in[2,11]\quad\text{and}\quad t\in[0,~{}0.6\times T]\$ (21)
@@
, where $\widetilde{.}$ denotes the known data. The first part, $\mathcal{L}_{\text{Sup}}$, is the supervised loss of the training where $\lambda=10$, which also includes the inlet loss. We used a limited number of velocity points (12 points) and 147 pressure points chosen randomly, using the total time snapshots, hence $t\in[0,T]$. The distribution of pressure and velocity points used for the supervised loss is given in Fig. 21. For the inlet function, we used the real parameters $(\widetilde{U}_{\text{max}},\widetilde{f})$ to calculate the real inlet velocity $\widetilde{\textbf{u}}_{\text{inlet}}$.
The second part, $\mathcal{L}_{\text{ICs}}$, represents the initial conditions. In this application, at $t=0$, velocity and pressure are set to 0, with $\eta=0.1$. We used $\mathcal{L}_{\text{BCs}}$ to impose the non-slip boundary condition for the wall and cylinder (annotated as $cyl$ in the equation). The physics loss $\mathcal{L}_{\text{PDE}}$ combines the residuals of the projections in the $x$ and $y$ directions for the conservation of momentum, $\mathcal{L}_{\text{eq1}}^{x}$ and $\mathcal{L}_{\text{eq1}}^{y}$, respectively, and the conservation of mass equation, $\mathcal{L}_{\text{eq2}}$ (Eqs. 10), which is used to calculate the velocity and pressure at positions where we do not have any supervised data. At the beginning of the training, $\lambda=0$, and we then increase it depending on the $\mathcal{L}_{\text{Sup}}$. The last part of the loss function, $\mathcal{L}_{\text{P}}$, represents the parameter loss of the inlet velocity function. For training, we used a batch size of 15 datasets. The validation dataset was used to study the convergence and select an optimal model. We used a similar training approach as Algo. 1. At the start of training, the learning rate is set to $lr=10^{-3}$, with $\alpha=10$, $\lambda=0$, and $\eta=0.1$. The learning rate $lr$ is decreased over epochs, with multiple thresholds adjusted to optimize training speed. After a pre-training phase where $\text{epoch}>\text{threshold}_{2}$, the physics loss is activated, and trainable variables are restricted to only the PINN model. The physics loss weight $\lambda$ is gradually increased when the supervised loss drops below a defined threshold.
@@ 5.3.3 Results and discussion
¯$\bar{\mathcal{E}}=\frac{1}{N_{t}}\sum_{t=1}^{N_{t}}\mathcal{E}(t_{i}),$ (24)
@@
Table 3: Aggregated model prediction error for $u_{x}$, $u_{y}$ and $p$, evaluated using : ¯$\bar{\mathcal{E}}_{\text{MAE}}$, ¯$\bar{\mathcal{E}}_{\text{median-AE}}$, and ¯$\bar{\mathcal{E}}_{\text{90\%-AE}}$. The table represent the distribution of these metrics across all test datasets, summarized by their mean, median (50%), and 90th percentile.
Variable Aggregation ¯$\bar{\mathcal{E}}_{\text{MAE}}$$\bar{\mathcal{E}}_{\text{median-AE}}$$\bar{\mathcal{E}}_{\text{90\%-AE}}$
@@
[…]
@@ 5.4.1 Samples
Table 4: Material properties of E-glass Fiber, PP (Isotactic Polypropylene), and PA66.
Property E-glass PP PA66 Units
@@ 5.4.2 Equipment
During the heating, the automate logs heating data w.r.t time (Fig. 28), namely temperature (in °$\textdegree C$) at the top and bottom also the oven air temperature, the % of heating power at the Top and at the Bottom. The acquisition frequency is set to 1 Hz, nevertheless in practice this frequency varies between 0.2 to 1 Hz. We also note that the quality of the pyrometer sensors is susceptible to some variation (for example, because of the color or the surface quality of the plates).
Figure 28: Examples of data collected during the heat process of two plates (PP at the top, and PA at the bottom). On the right, we have the temperature at the top and bottom during the heating process, and on the left, we have the power of the heating at the top and bottom. As seen, we used a special protocol, where we set the power at the bottom to 0% at a random instant $T_{pdb}$.
@@ 5.4.4 Numerical data generation
$16.3-18.3$ (high)
$L$ (mm) 1.58 – 2.41 Linearly increasing.
@@
For each dataset (combination of the variations in Tab. 5), we simulate the 1D heating problem (Eq. 11) using the backward Euler scheme, where $h=8$ is the natural thermal convection coefficient. We study the convergence of the numerical solution by decreasing both the time step $\Delta t$ and the spatial step $\Delta x$. For each dataset, we generate both cases: one where there is a drop in the power heating percentage at the bottom ($x=L$), and one without the drop in power (Fig. 30). This means that the model is trained in both cases. We did this to give the model the ability to learn the impact of the drop in heat power percentage. For each simulation, we use $t\in[0,~{}T]$, where $T=45~{}\text{s}$, while for the ICs, as given in Tab. 5, for each dataset, we generate $T(t=0,~{}x=0)$ and $T(t=0,~{}x=L)$, and then we calculate the ICs $T(t=0,~{}x)$ for each $x\in]0,L[$ using an affine function based on $T(x=0,t)$ and $T(x=L,t)$.
@@ 5.4.5 Model architecture and training process
The architecture used in this study is shown in Fig. 31. In this application, the Deep Set takes as input $X_{DS}=[time,~{}T_{0},~{}T_{L},~{}P_{0},~{}P_{L},~{}L]$, where time represents the time of each data snapshot. The model is trained with a variable $\Delta T$, which reflects the fact that, in real data, the frequency of acquisition is not constant. We also provide the corresponding temperatures $T_{0}=T(t_{i},x=0)$ and $T_{L}=T(t_{i},x=L)$ for each time snapshot, as well as the power inputs $P_{0}=P(t_{i},x=0)$ and $P_{L}=P(t_{i},x=L)$, expressed as percentages (not as heat flux), along with the constant thickness $L$. In this application, we used Random Fourier Features [29] as a projection of the QPs ($time$ and $x$), simply as an encoding step. Since the QPs are identical across all batches during training, the projection is computed once for a single batch and then repeated for the entire batch dimension. This approach helps to reduce redundant computations. This approach helps with the convergence of the model. We used only a PINN, without any identification blocks, as the range of diffusivity is relatively narrow, making the inclusion of additional blocks unnecessary for this application.
@@
, where the first part, $\mathcal{L}_{\text{PDE}}$, represents the physics loss (as given in Eq. 11), the second part, $\mathcal{L}_{\text{IC}}$, represents the initial condition loss, and the third part, $\mathcal{L}_{\text{BC}}$, represents the boundary condition loss (as given in Eq. 11). The term $\mathcal{L}_{\text{SUP}}$ represents the supervised loss, which is simply the error between $T_{\text{real}}(t,x)$ and $T_{\text{pred}}(t,x)$, for $x\in\{0,L\}$ and $t\in[0,T]$. This error varies during the training, both before and after the drop of heat power $P_{x=L}$ to 00 (see Fig. 30 and 32). At the beginning of training, the supervised loss is prioritized by initializing the weights as $\beta_{1}=0$, $\beta_{2}=5$, $\beta_{3}=1$, and $\beta_{4}=10$. As training progresses, the physics loss weight $\beta_{1}$ is gradually increased once both the supervised and initial condition losses are below specific thresholds. Additionally, the learning rate is adaptively decayed based on the training epochs to ensure stable convergence.
@@
Deep Set $\text{gelu}(X)$ 3 64 32 Aggregation is performed using: |.$\sum$|.|∑ | . |.

### v2

@@ 1 Introduction
The rapid growth of research in physics-informed machine learning [3] (PIML) has led to innovative methods that integrate physics-based knowledge into data-driven models [4, 5, 6]. These approaches benefit from advancements in both machine learning algorithms and hardware acceleration, making them increasingly accessible. Unlike conventional neural networks, physics-informed models incorporate governing physical laws into the learning process, improving their ability to generalize beyond available supervised data. A notable example is Physics-Informed Neural Networks (PINNs), which leverage both physics-based constraints and observed data to solve Partial and Ordinary Differential Equations (PDEs/ODEs). PINNs have gained considerable interest due to their successful applications in direct problems, in various fields such fluid dynamics [7], multiphysics coupled problems [8, 9], as well as their ability to handle multi-fidelity data sources [10]. PINNs could also be used for inverse problems [11, 12, 13]. They have also been benchmarked against Finite Element Methods (FEM) [14], demonstrating very good performance. Furthermore, PINNs have shown robustness in solving inverse problems, effectively handling noisy data to infer accurate underlying dynamics. This ability to explore physics through data-driven methods has become a highly active research area.
@@ 2 Identification methods and PINN
$\underset{\xi}{\text{argmin}}\quad\lambda\mathcal{L}_{\text{BC}}(\xi)+\mathcal{L}_{\text{IC/ SUP}}(\xi)+\beta\mathcal{L}_{\text{PDE}}(\xi)$ (5)
@@ 4 Architecture of the combination of Sequence Encoder and Physics Informer Neural Networks
Figure 6: Overview of the PINN-SE model that combines two Sequences Encoders (Deep Sets or Set Transformer) with PINN and identification blocks, along with the training logic of the model. In this example architecture, we use two Sequence Encoders: the first to encode sensor data $s_{i}$, and the second to encode boundary conditions (BCsBCs). Each SE generates a feature vector, which are then concatenated into a global feature vector $\mathbf{F}_{\text{vect}}$ and used by both the PINN and the identification blocks.
@@ 5.2.1 Data generation
[…]
@@ 5.2.2 Model architecture and training
The model is trained using the expanding window method (Example in Fig. 10). This approach leverages the advantages of Deep sets, that can encode variable sequence lengths and map them to dynamic behaviors. This means the model is trained to integrate new time snapshot data to update its predictions. In this application, the model is trained using 362 datasets selected randomly (using Sobol sampling), the rest is used for validation and testing. The model architecture is given in Fig. 9. The model input, $X_{DS}$, represents time and corresponding state variable snapshots. We use a time step of $\Delta T_{DS}=0.5~s$ for the Deep sets input. The input tensor is $X_{DS}\in\mathbb{R}^{b\times N\times d}$, where $b$ is the batch size, $N\in[10,40]$ is the number of snapshot, and $d=4$ is the input dimension (time, $x$, $y$, and $z$). This input is projected to a higher-dimensional representation $\Phi\in\mathbb{R}^{b\times N\times d_{f.v}}$, where $d_{f.v}$ is the dimension of the feature vector. We then apply an aggregation function to compute the feature vector $\mathbf{F}_{\text{vect}}\in\mathbb{R}^{b\times 1\times d_{f.v}}$. The choice of aggregation function is crucial and must be made carefully to ensure generalization. In this application, we used the following aggregation Eq. 13:
$\left\|.\right\|_{2}^{2}\equiv\sum_{i=1}^{N}\frac{\Phi_{i}^{2}}{N}\quad,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{where}}\quad{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Phi=\left[\Phi_{1}...\Phi_{i}...\Phi_{N}\right]}\quad{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Phi_{i}\in\mathbb{R}^{b\times 1\times d_{f.v}}}$ (13)
@@
Figure 9: The PINN-SE model used for the first application (Rössler dynamics) operates as follows: the Deep set takes as input the timetime and state variables $x$, $y$, and $z$, given as $X_{DS}\in\mathbb{R}^{\text{b}\times N\times 4}$, and predicts feature vector $\mathbf{F}_{\text{vect}}\in\mathbb{R}^{b\times 1\times d_{f.v}}$. This feature vector is then used directly by the identification blocks (MLPs) to predict the parameter $c$. The feature vector is also concatenated with the time snapshots for which we want to predict the state variables $x$, $y$, and $z$ using the PINN blocks. The dynamics loss (Eq. 15) is calculated using automatic differentiation.
@@
Where $\mathcal{L}_{\text{sup}}$ is the supervised loss, and I(t)=[x(t),y(t),z(t)]I(t)=[x(t),~y(t),~z(t)] represents the labeled noisy points used for supervision loss. We employed sparse input sampling with $\Delta T_{\text{sup}}=0.5$ for the supervised loss. The model prediction, $\widetilde{I(t)}$, is computed as shown in Eq. 16:
@@
Deep set 80 4 $\text{gelu}(X)$ -
Pooling - $\left\|.\right\|_{2}^{2}$
Post-Pooling 80 3 $\text{gelu}(X)$ -
@@
Identification 80 3 $\text{gelu}(X)$ -
@@
12:  if Validation error of parameters << threshold then
@@ 5.2.3 Results and analysis
[…]
@@ 5.3.1 Data generation and preparation
$\mathbf{u}$ $=[0,0]\quad\text{on the walls and obstacle (no-slip condition)},$
@@ 5.3.2 Model architecture and training process
$\textbf{X}_{DS}=\{t,~p_{1},~\cdots,~p_{i},~\cdots,~p_{12}\}\in\mathbb{R}^{b\times N_{DS}\times 13}\quad i\in[2,11]\quad\text{and}\quad t\in[0,~0.6\times T]\\$ (21)
@@
[…]
The second part, $\mathcal{L}_{\text{ICs}}$, represents the initial conditions. In this application, at $t=0$, velocity and pressure are set to 0, with $\eta=0.1$. We used $\mathcal{L}_{\text{BCs}}$ to impose the non-slip boundary condition for the wall and cylinder (annotated as cylcyl in the equation). The physics loss $\mathcal{L}_{\text{PDE}}$ combines the residuals of the projections in the $x$ and $y$ directions for the conservation of momentum, $\mathcal{L}_{\text{eq1}}^{x}$ and $\mathcal{L}_{\text{eq1}}^{y}$, respectively, and the conservation of mass equation, $\mathcal{L}_{\text{eq2}}$ (Eqs. 10), which is used to calculate the velocity and pressure at positions where we do not have any supervised data. At the beginning of the training, $\lambda=0$, and we then increase it depending on the $\mathcal{L}_{\text{Sup}}$. The last part of the loss function, $\mathcal{L}_{\text{P}}$, represents the parameter loss of the inlet velocity function. For training, we used a batch size of 15 datasets. The validation dataset was used to study the convergence and select an optimal model. We used a similar training approach as Algo. 1. At the start of training, the learning rate is set to $lr=10^{-3}$, with $\alpha=10$, $\lambda=0$, and $\eta=0.1$. The learning rate $lr$ is decreased over epochs, with multiple thresholds adjusted to optimize training speed. After a pre-training phase where $\text{epoch}>\text{threshold}_{2}$, the physics loss is activated, and trainable variables are restricted to only the PINN model. The physics loss weight $\lambda$ is gradually increased when the supervised loss drops below a defined threshold.
@@ 5.3.3 Results and discussion
$\bar{\mathcal{E}}=\frac{1}{N_{t}}\sum_{t=1}^{N_{t}}\mathcal{E}(t_{i}),$ (24)
@@
Table 3: Aggregated model prediction error for $u_{x}$, $u_{y}$ and $p$, evaluated using : $\bar{\mathcal{E}}_{\text{MAE}}$, $\bar{\mathcal{E}}_{\text{median-AE}}$, and $\bar{\mathcal{E}}_{\text{90\%-AE}}$. The table represent the distribution of these metrics across all test datasets, summarized by their mean, median (50%), and 90th percentile.
Variable Aggregation $\bar{\mathcal{E}}_{\text{MAE}}$ $\bar{\mathcal{E}}_{\text{median-AE}}$ $\bar{\mathcal{E}}_{\text{90\%-AE}}$
@@
$\text{Scaled-MAE}_{\text{model}}=\frac{1}{\max(\widetilde{u}_{x}^{\text{inlet}})-\min(\widetilde{u}_{x}^{\text{inlet}})}\cdot\sum\left|u_{x}^{\text{inlet}}-\widetilde{u}_{x}^{\text{inlet}}\right|,$ (25)
$\text{Scaled-MAE}_{\text{comparison}}=\frac{1}{\max(\widetilde{u}_{x}^{\text{inlet}})-\min(\widetilde{u}_{x}^{\text{inlet}})}\cdot\sum\left|\widehat{u}_{x}^{\text{inlet}}-\widetilde{u}_{x}^{\text{inlet}}\right|$ (26)
@@ 5.4.1 Samples
Table 4: Material properties of E-glass Fiber, $P$ (Isotactic Polypropylene), and PA66.
Property E-glass $P$ PA66 Units
@@ 5.4.2 Equipment
During the heating, the automate logs heating data w.r.t time (Fig. 28), namely temperature (in $\textdegree C$) at the top and bottom also the oven air temperature, the % of heating power at the Top and at the Bottom. The acquisition frequency is set to 1 Hz, nevertheless in practice this frequency varies between 0.2 to 1 Hz. We also note that the quality of the pyrometer sensors is susceptible to some variation (for example, because of the color or the surface quality of the plates).
Figure 28: Examples of data collected during the heat process of two plates ($P$ at the top, and PA at the bottom). On the right, we have the temperature at the top and bottom during the heating process, and on the left, we have the power of the heating at the top and bottom. As seen, we used a special protocol, where we set the power at the bottom to 0% at a random instant $T_{pdb}$.
@@ 5.4.4 Numerical data generation
16.318.316.3-18.3 (high)
$L$ ($m$) 1.58 – 2.41 Linearly increasing.
@@
For each dataset (combination of the variations in Tab. 5), we simulate the 1D heating problem (Eq. 11) using the backward Euler scheme, where $h=8$ is the natural thermal convection coefficient. We study the convergence of the numerical solution by decreasing both the time step $\Delta t$ and the spatial step $\Delta x$. For each dataset, we generate both cases: one where there is a drop in the power heating percentage at the bottom ($x=L$), and one without the drop in power (Fig. 30). This means that the model is trained in both cases. We did this to give the model the ability to learn the impact of the drop in heat power percentage. For each simulation, we use $t\in[0,~T]$, where $T=45~\text{s}$, while for the ICs, as given in Tab. 5, for each dataset, we generate T(t=0,x=0)T(t=0,~x=0) and T(t=0,x=L)T(t=0,~x=L), and then we calculate the ICs T(t=0,x)T(t=0,~x) for each $x\in]0,L[$ using an affine function based on $T(x=0,t)$ and $T(x=L,t)$.
@@ 5.4.5 Model architecture and training process
The architecture used in this study is shown in Fig. 31. In this application, the Deep Set takes as input $X_{DS}=[time,~T_{0},~T_{L},~P_{0},~P_{L},~L]$, where time represents the time of each data snapshot. The model is trained with a variable $\Delta T$, which reflects the fact that, in real data, the frequency of acquisition is not constant. We also provide the corresponding temperatures $T_{0}=T(t_{i},x=0)$ and $T_{L}=T(t_{i},x=L)$ for each time snapshot, as well as the power inputs $P_{0}=P(t_{i},x=0)$ and $P_{L}=P(t_{i},x=L)$, expressed as percentages (not as heat flux), along with the constant thickness $L$. In this application, we used Random Fourier Features [38] as a projection of the QPs (timetime and $x$), simply as an encoding step. Since the QPs are identical across all batches during training, the projection is computed once for a single batch and then repeated for the entire batch dimension. This approach helps to reduce redundant computations. This approach helps with the convergence of the model. We used only a PINN, without any identification blocks, as the range of diffusivity is relatively narrow, making the inclusion of additional blocks unnecessary for this application.
@@
, where the first part, $\mathcal{L}_{\text{PDE}}$, represents the physics loss (as given in Eq. 11), the second part, $\mathcal{L}_{\text{IC}}$, represents the initial condition loss, and the third part, $\mathcal{L}_{\text{BC}}$, represents the boundary condition loss (as given in Eq. 11). The term $\mathcal{L}_{\text{SUP}}$ represents the supervised loss, which is simply the error between $T_{\text{real}}(t,x)$ and $T_{\text{pred}}(t,x)$, for $x\in\{0,L\}$ and $t\in[0,T]$. This error varies during the training, both before and after the drop of heat power $P_{x=L}$ to 0 (see Fig. 30 and 32). At the beginning of training, the supervised loss is prioritized by initializing the weights as $\beta_{1}=0$, $\beta_{2}=5$, $\beta_{3}=1$, and $\beta_{4}=10$. As training progresses, the physics loss weight $\beta_{1}$ is gradually increased once both the supervised and initial condition losses are below specific thresholds. Additionally, the learning rate is adaptively decayed based on the training epochs to ensure stable convergence.
@@
Deep Set $\text{gelu}(X)$ 3 64 32 Aggregation is performed using: $\sum|.|$.